# Faster and Better 3D Splatting via Group Training

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for novel view synthesis, demonstrating remarkable capability in high-fidelity scene reconstruction through its Gaussian primitive representations. However, the computational overhead induced by the massive number of primitives poses a significant bottleneck to training efficiency. To overcome this challenge, we propose Group Training, a simple yet effective strategy that organizes Gaussian primitives into manageable groups, optimizing training efficiency and improving rendering quality. This approach shows universal compatibility with existing 3DGS frameworks, including vanilla 3DGS and Mip-Splatting, consistently achieving accelerated training while maintaining superior synthesis quality. Extensive experiments reveal that our straightforward Group Training strategy achieves up to 30% faster convergence and improved rendering quality across diverse scenarios.
@@ 1 Introduction
Our code will be made publicly available upon acceptance of this paper ( but is now available in the supplementary materials for reviewing).
@@ 2 Related Work
Recent studies have driven significant advancements in 3DGS across several areas, notably improving rendering quality [2, 30], memory efficiency [24, 5], and scalability for large scenes [15].
Beyond these enhancements, additional research has focused on increasing the efficiency of 3DGS optimizers [10] and refining Gaussian Densification [18, 17] and Pruning strategies [5, 6, 25]. These improvements yield more compact 3DGS models that support faster reconstruction without sacrificing detail.
@@ 3 Methodology
In Sec. 3.1, we will provide a comprehensive introduction to the Group Training method and its multifunctional advantages in 3D Gaussian Splatting (3DGS). Section 3.2 introduces two sampling paradigms: Random Sampling and Prioritized Sampling (incorporating importance score, opacity-based, and volume-based criteria), with comparative analysis of their computational efficiency and performance trade-offs. We mathematically demonstrate that Opacity-based Prioritized Sampling achieves effective 3DGS densification and efficient rendering. Finally, in Sec. 3.3 , we summary training characteristics of 3DGS with Group Training, and the Sampling Strategies are identified.
@@ 3.1 Group Training
Before the next grouping, the Cached Group is merged with the Under-training Group, after which the new and groups are resampled. A visual representation of the Group Training process is provided in Fig. 3 This procedure ensures that each Gaussian primitive has an opportunity to contribute to the training process, and effectively reducing the number of Under-training Gaussians while retaining those deemed “important.”
@@ 3.2 Effective and Efficient Sampling Strategies
Within each groupinsg, a subset of Gaussians is sampled from the merged Gaussians. The simplest sampling strategy is Random Sampling (RS). We propose attribute-driven Prioritized Sampling to optimize 3DGS further.
We evaluated various attributes , including opacity and volume-based approaches [5] and importance scores [25, 6, 32, 8] (see appendix for detailed comparisons). Among these, Opacity-based Prioritized Sampling demonstrated superior performance in both training speed and rendering quality. As mathematically proven in Sec. 3.2.1 and 3.2.1, opacity intrinsically governs two critical 3DGS properties, both experimentally validated.
@@ 3.2.1 Sampling Strategy for Better Densification
For any Gaussian primitive projected onto the imaging plane with 2D center coordinates . The partial derivatives of loss w.r.t the projected coordinates are , which govern the densification criterion by:
@@ 3.2.2 Sampling Strategy for Faster Rendering
where the parameter denotes the number of Gaussians required to render pixel . As established in 3DGS [12], the determination of hinges on achieving opacity saturation within a pixel’s accumulated value (Transmittance ). This critical threshold implementation introduces a fundamental constraint on rendering efficiency: during forward rendering operations, the sequential traversal of Gaussian primitives inherently prevents parallel computation of values. Consequently, emerges as the principal determinant of rendering speed, where any reduction on directly translates to accelerated renderings.
@@ 4.1 Experimental Settings
Implementation Details. The Group Training method was integrated as a plug-in to facilitate its utilisation in conjunction with 3DGS. We implemented our method within both 3D Gaussian Splatting [12]111https://github.com/graphdeco-inria/gaussian-splatting and Mip-Splatting [30]222https://github.com/autonomousvision/mip-splatting.
@@
Dataset and Metrics. The efficacy of our method is evaluated on the Mip-NeRF360 [1]333https://jonbarron.info/mipnerf360/, Tanks & Temples [13]444https://www.tanksandtemples.org/, Deep Blending [9]555http://visual.cs.ucl.ac.uk/pubs/deepblending/, and NeRF-Synthetic [20]666https://www.matthewtancik.com/nerf using a single RTX 3090. Subsequently, the model’s reconstruction performance is evaluated using PSNR, SSIM, and LPIPS. Furthermore, the training time, peak memory usage, and model size for each scene are recorded.
@@ 4.5 Discussion
Why faser? Group Training significantly accelerates reconstruction primarily by reducing the number of Under-training Gaussian primitives, thereby decreasing the rendering burden on the model. This speedup is observed even when the final model size is larger than the baseline. In alignment with our earlier analysis in Sec. 3.2, the OPS also contributes notably to this speedup effect. During the densification phase, Group Training with OPS preserves high-opacity Gaussians within the Under-Training Group, without which could lead to instances of both under-reconstruction and over-reconstruction [12, 29]. These issues can lead to redundant densification. As a result, models reconstructed with Group Training and OPS contain fewer redundant Gaussians, leading to smaller model sizes and faster reconstruction overall. Additionally, OPS prioritizes sampling high-opacity Gaussians, which rapidly drives saturation, which also discussed in Sec. 3.2. Focusing on these most impactful Gaussians reduces the number of Gaussians needed for splatting, further enhancing training and rendering efficiency.
Why better? Group Training with OPS prioritizes the retention and optimization of high-opacity Gaussian primitives, ensuring that elements most crucial to scene rendering are refined more effectively. The Gaussian primitives with higher opacity inherently have a greater influence on the final rendered image [5, 26, 18, 21], so emphasizing their optimization leads to significant improvements in overall rendering quality. Besides, cyclic resampling temporarily excluding a randomly selected subset of Gaussian primitives, each remaining primitive is compelled to learn how to independently represent essential scene features leading to a more robust representation of the scene.
@@ Appendix B Efficiency And Effectiveness For Various Sampling Strategies
where represents the sampling metrics (opacity [5], volume [15] or importance score 777Based on code: https://github.com/fatPeter/mini-splatting.git [25, 6, 32]) of Gaussian primitive , and is the total number of Gaussian primitives. We also evaluated the metric which both Opacity and Volume are considered simultaneously, referred to as the Volume & Opacity-based method, as applied in [5]. The sampling metric for Volume & Opacity-based Prioritized Sampling is computed as follows:
@@
Appendix C Methodological Applicability
@@
Appendix D Temporal Evolution of Under-Training Gaussian Primitives.
Appendix E Detailed Experimental Results for All Scenes

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for novel view synthesis, demonstrating remarkable capability in high-fidelity scene reconstruction through its Gaussian primitive representations. However, the computational overhead induced by the massive number of primitives poses a significant bottleneck to training efficiency. To overcome this challenge, we propose Group Training, a simple yet effective strategy that organizes Gaussian primitives into manageable groups, optimizing training efficiency and improving rendering quality. This approach shows universal compatibility with existing 3DGS frameworks, including vanilla 3DGS and Mip-Splatting, consistently achieving accelerated training while maintaining superior synthesis quality. Extensive experiments reveal that our straightforward Group Training strategy achieves up to 30% faster convergence and improved rendering quality across diverse scenarios. Project Website: https://chengbo-wang.github.io/3DGS-with-Group-Training/.
@@ 2 Related Work
Recent studies have driven significant advancements in 3DGS across several areas, notably improving rendering quality [2, 33, 13], memory efficiency [25, 5, 27], and scalability for large scenes [16].
Beyond these enhancements, additional research has focused on increasing the efficiency of 3DGS optimizers [10] and refining Gaussian Densification [19, 18, 31] and Pruning strategies [5, 6, 26]. These improvements yield more compact 3DGS models that support faster reconstruction but sacrificing detail.
@@ 3 Methodology
In Sec. 3.1, we will provide a comprehensive introduction to the Group Training method and its multifunctional advantages in 3D Gaussian Splatting (3DGS). Section 3.2 introduces two sampling paradigms: Random Sampling and Prioritized Sampling (incorporating importance score, opacity-based, and volume-based criteria), with comparative analysis of their computational efficiency and performance trade-offs. We mathematically demonstrate that Opacity-based Prioritized Sampling achieves effective 3DGS densification and efficient rendering. Finally, in Sec. 3.3 , we summarize training characteristics of 3DGS with Group Training, and the Sampling Strategies are identified.
@@ 3.1 Group Training
Before the next grouping, the Cached Group is merged with the Under-training Group, after which the new and groups are resampled. A visual representation of the Group Training process is provided in Fig. 3. This procedure ensures that each Gaussian primitive has an opportunity to contribute to the training process, and effectively reducing the number of Under-training Gaussians while retaining those deemed “important.”
@@ 3.2 Effective and Efficient Sampling Strategies
Within each grouping, a subset of Gaussians is sampled from the merged Gaussians. The simplest sampling strategy is Random Sampling (RS). We propose attribute-driven Prioritized Sampling to optimize 3DGS further.
We evaluated various attributes, including opacity and volume-based approaches [5] and importance scores [26, 6, 35, 8] (see appendix for detailed comparisons). Among these, Opacity-based Prioritized Sampling demonstrated superior performance in both training speed and rendering quality. As mathematically proven in Sec. 3.2.1 and 3.2.2, opacity intrinsically governs two critical 3DGS properties, both experimentally validated.
@@ 3.2.1 Sampling Strategy for Better Densification
For any Gaussian primitive projected onto the imaging plane with 2D center coordinates , the partial derivatives of loss w.r.t the projected coordinates are , which govern the densification criterion by:
@@ 3.2.2 Sampling Strategy for Faster Rendering
As established in 3DGS [12], the determination of hinges on achieving opacity saturation within a pixel’s accumulated value (Transmittance ). This critical threshold implementation introduces a fundamental constraint on rendering efficiency: during forward rendering operations, the sequential traversal of Gaussian primitives inherently prevents parallel computation of values. Consequently, emerges as the principal determinant of rendering speed, where any reduction on directly translates to accelerated renderings.
@@ 4.1 Experimental Settings
Implementation Details. The Group Training method was integrated as a plug-in to facilitate its utilisation in conjunction with 3DGS. We implemented our method within both 3D Gaussian Splatting [12]††https://github.com/graphdeco-inria/gaussian-splatting and Mip-Splatting [33]††https://github.com/autonomousvision/mip-splatting.
@@
Dataset and Metrics. The efficacy of our method is evaluated on the Mip-NeRF360 [1]††https://jonbarron.info/mipnerf360/, Tanks & Temples [14]††https://www.tanksandtemples.org/, Deep Blending [9]††http://visual.cs.ucl.ac.uk/pubs/deepblending/, and NeRF-Synthetic [21]††https://www.matthewtancik.com/nerf using a single RTX 3090. Subsequently, the model’s reconstruction performance is evaluated using PSNR, SSIM, and LPIPS. Furthermore, the training time, peak memory usage, and model size for each scene are recorded.
@@ 4.5 Discussion
Why faster? Group Training significantly accelerates reconstruction primarily by reducing the number of Under-training Gaussian primitives, thereby decreasing the rendering burden on the model. This speedup is observed even when the final model size is larger than the baseline. In alignment with our earlier analysis in Sec. 3.2, the OPS also contributes notably to this speedup effect. During the densification phase, Group Training with OPS preserves high-opacity Gaussians within the Under-Training Group, without which could lead to instances of both under-reconstruction and over-reconstruction [12, 32]. These issues can lead to redundant densification. As a result, models reconstructed with Group Training and OPS contain fewer redundant Gaussians, leading to smaller model sizes and faster reconstruction overall. Additionally, OPS prioritizes sampling high-opacity Gaussians, which rapidly drives saturation, which also discussed in Sec. 3.2. Focusing on these most impactful Gaussians reduces the number of Gaussians needed for splatting, further enhancing training and rendering efficiency.
Why better? Group Training with the OPS prioritizes the retention and optimization of high-opacity Gaussian primitives. This achieves comparable or superior fidelity to RS, while significantly reducing the model size. The Gaussian primitives with higher opacity inherently have a greater influence on the final rendered image [5, 28, 19, 22], so emphasizing their optimization leads to significant improvements in overall rendering quality. Besides, cyclic resampling temporarily excluding a randomly selected subset of Gaussian primitives, each remaining primitive is compelled to learn how to independently represent essential scene features leading to a more robust representation.
@@ 5 Conclusion
Acknowledgements
This work is supported by the National Key R&D Program of China (No.2022ZD0119000), Hunan Provincial Key R&D Program of China (No.2024JK2020 and 2024JK2021), Hunan Provincial Natural Science Foundation of China (No.2024JJ10027), Young Talents of Huxiang (No.Z202433000575), and Changsha Science Fund for Distinguished Young Scholars (kq2306002).
Supplementary Material
@@ Appendix B Efficiency And Effectiveness For Various Sampling Strategies
where represents the sampling metrics (opacity [5], volume [16] or importance score ††Based on code: https://github.com/fatPeter/mini-splatting.git [26, 6, 35]) of Gaussian primitive , and is the total number of Gaussian primitives. We also evaluated the metric which both Opacity and Volume are considered simultaneously, referred to as the Volume & Opacity-based method, as applied in [5]. The sampling metric for Volume & Opacity-based Prioritized Sampling is computed as follows:
@@
Appendix C Temporal Evolution of Under-Training Gaussian Primitives
Appendix D Comparison of Scene Representation Efficiency
Fig. 9 presents a comparative analysis for the ”train” scene reconstruction using 3D Gaussian Splatting (3DGS) under baseline conditions versus Group Training. It compares the per-iteration FPS and the number of hitten Gaussians. And the baseline method required 12.5 minutes to reach a PSNR of 21.985 dB, whereas Group Training with OPS acceleration attained a PSNR of 22.156 dB in just 9.3 minutes. These measurements confirm that Group Training consistently accelerates rendering with a substantial reduction in hitten Gaussians count during training. Consequently, it demonstrates higher scene representation efficiency by utilizing fewer Gaussian primitives without compromising reconstruction quality.
Appendix E Distribution of Gaussian Primitives in Imaging Plane Space
Fig. 10 illustrates the projection of rendering Gaussian primitives onto the imaging plane. Our Group Training approach significantly reduces the number of primitives required per image compared to the baseline, without compromising rendering quality, and further improves reconstruction speed.
Appendix F Methodological Applicability
@@
Appendix G Detailed Experimental Results for All Scenes